# Excitonic Theory of the Ultrafast Optical Response of 2D-Quantum-Confined Semiconductors at Elevated Densities

### v1

@@ IV Excitonic Rabi Oscillations
in units of one over area. Within an expansion in electron-hole operator pairs Katsch et al. (2018); Ivanov and Haug (1993), cf. Eq. (7), the total electron density $N$ can also be expressed as:
@@
To study the exciton-density dynamics in Eq. (46), consider two cases, a monolayer MoSe2 (strong Coulomb coupling) and a GaAs QW (less strong Coulomb coupling). Both systems are resonantly excited at $\hbar\omega_{\text{P}}=E_{1s}$ using a circularly polarized 2-ps pulse. We solve the excitonic equations of motion for the excitonic transitions $P_{1s}$ and excitonic occupations $N_{1s,\mathbf{Q}}$. The equations of motion for the two-exciton continua $B_{+,\zeta,1s}$ and exciton-two-exciton continua $Z_{+,\zeta,1s,\mathbf{Q}}$ for $\xi=\xi^{\prime}=K,\uparrow$ (monolayer MoSe2) or $\xi=\xi^{\prime}=\text{heavy-hole~exciton}$ (GaAs QW) are eliminated via a Markov approximation, i.e., we assume an instantaneous build-up of exciton-exciton interaction and neglect memory effects Axt et al. (2004), which is reasonable given the comparably long pulse duration. We take only $\mu=1s$ excitonic transitions and occupations into account and neglect any intraexcitonic coherences $N_{\mu,\nu\neq\mu,\mathbf{Q}}$, but include $1s$, $2s$ and $3s$ excitonic states in the calculation of the symmetric two-exciton continua $B_{+,\zeta}$ and symmetric exciton-two-exciton continua $Z_{+,\zeta}$. Due to the circular excitation, antisymmetric ”-”-configurations, which host a bound biexciton/exciton-biexciton, are not induced.
For comparison with the excitonic theory, we also solve the well-established semiconductor Bloch equations in Hartree-Fock limit Lindberg and Koch (1988) to obtain the density of uncorrelated electrons in Eq. (45):
@@
All material parameters are listed in Tab. 2. To ensure comparable conditions, we assume equal nonradiative dephasings $\gamma_{\text{nrad}}$ for the MoSe2 monolayer and GaAs QW and neglect any contributions due to reradiation Knorr et al. (1996). In a MoSe2 monolayer, a total (half) linewidth of 1 meV corresponds roughly to the radiatively-limited case, as the radiative (half) linewidth already approaches 1 meV, while in a GaAs QW, a total (half) linewidth of 1 meV is much larger than the radiative linewidth of approximately tens of µeV Schäfer et al. (2025).
@@
Effective well width $d_{\text{eff}}$ $1.3d$ Liu (1994)
@@
Static dielectric constant of GaAs 12.46 (5 K) Moore and Holm (1996)
@@
h-BN-encapsulated MoSe2 monolayer
Monolayer width $d$ 0.6527 nm Kylänpää and Komsa (2015)
@@
Static dielectric constant of bulk MoSe2 $\sqrt{8.2\cdot 17.7}$ Laturia et al. (2018)
@@
Plasmon energy of MoSe2 $\hbar\omega_{\text{pl}}$ 22 eV Kumar and Ahluwalia (2012)
@@
CMR Andersen et al. (2015))
Static dielectric constant of bulk h-BN 4.8 Latini et al. (2015)
Interlayer gap $h$ 0.3 nm Florian et al. (2018); Drüppel et al. (2017)
@@
GaAs quantum well MoSe2 monolayer
@@
Figure 7: Dynamics of the total electron density NN from Eq. (45) as solution of the semiconductor Bloch equations (SBE) (a,f) compared to the total exciton density NN from Eq. (46) as a solution of the excitonic Bloch equations in the coherent limit (COH) (b,g), with additional exciton-exciton scattering contributions (COH+\,+\,$X$) (c,h) and with additional correlated contributions (COH+\,+\,CORR) (d,i) and with all contributions (COH+\,+\,CORR+\,+\,$X$) (e,j) at various pulse areas $\Theta$, cf. Eq. (26), for a 8-nm GaAs quantum well (left column) and for a MoSe2 monolayer (right column). The grey-shaded area depicts the normalized pulse intensity.
@@
In the following, we dissect the corresponding interaction contributions to the exciton dynamics and compare the emerging dynamics of a MoSe2 ML and a GaAs QW for circularly-polarized excitation (GaAs QW vs. MoSe2 monolayer). Then, we identify the role of the coherent and (optically dark) incoherent exciton density (Coherent vs. incoherent exciton density) and examine the momentum distribution of optically dark excitons (Momentum distribution of optically dark excitons). At last, we compare circular and linear excitation in a MoSe2 ML (Circular vs. linear excitation).
@@
GaAs QW vs. MoSe2 monolayer: In Fig. 7, we depict the dynamics of the electron density, cf. Eq. (45), by solving the semiconductor Bloch equations (SBEs) in Eq. (47) and the dynamics of the total excitonic density, cf. Eq. (46), as solutions of the full excitonic Bloch equations in Eqs. (21)–(44) within the limits discussed above at various pulse areas, defined in Eq. (26). The left column depicts simulations for a GaAs QW and the right column depicts the simulations for a h-BN-encapsulated MoSe2 monolayer.
@@
In Fig. 7(a,f), which shows the results from solving the SBEs, Eq. (47), Rabi flops are observed, whose period increases with increasing pulse area. In Fig. 7(a) (GaAs QW), up to four Rabi flops can be observed at a pulse area of $2\pi$, while Fig. 7(f) (MoSe2 ML) shows a multitude of Rabi flops. As Rabi oscillations are enhanced due to the occurrence of Coulomb-induced internal fields, cf. last term in the first line in Eq. (47), and the Coulomb interaction in a MoSe2 ML is much stronger compared to a GaAs QW due to the reduced screening, more Rabi flops occur in a MoSe2 monolayer.
@@
In Fig. 7(b,g), we depict the coherent (COH) limit of the excitonic Bloch equations, cf. Eq. (21) and Eq. (30), which corresponds to the excitonic description in Ref. Knorr et al. (1994), which does not include nonradiative dephasing. Here, the exciton density decays within 3 ps, as no long-lived incoherent occupations $N_{1s,\mathbf{Q}}$ are induced. In the GaAs QW, Fig. 7(b), one Rabi flop occurs at the largest pulse area, as the nonradiative dephasing $\hbar\gamma_{\text{nrad}}$ suppresses them. In a MoSe2 monolayer, Fig. 7(g), Rabi flops occur in a similar number compared to the SBE-case in Fig. 7(f), since the stronger Coulomb-induced field renormalization counterbalances the nonradiative dephasing on short timescales.
@@
In Fig. 7(c,h), we depict the coherent limit with exciton-exciton interaction corresponding to two-exciton continua (COH+\,+\,$X$), cf. Eq. (21), Eq. (30), Eq. (32) and Eq. (29), which corresponds to the model developed in Refs. Katsch et al. (2019, 2020); Katsch and Knorr (2020). In this limit, the two-exciton-induced dephasing, which corresponds to the coupling of excitonic transitions to two-exciton continua (excitation-induced dephasing) totally suppresses the (coherent) Rabi oscillations in the GaAs QW and greatly suppresses them in a MoSe2 monolayer, but not completely, as two Rabi flops of small amplitude still remain. The overall induced coherent exciton density is larger compared to Fig. 7(b,g), as the exciton-exciton interaction reduces the overall excitation-dependent blue shift due to bandgap reduction and phase-space filling in Eq. (30) (first line) enabling a more efficient optical excitation of electron-hole pairs. This is especially relevant in a MoSe2 ML, as the enhanced Coulomb interaction causes a stronger exciton-exciton interaction and hence a stronger excitation-induced dephasing and -attenuation of the blue shift.
@@
In Fig. 7(d,i), we turn off the exciton-exciton (XX) interaction in Eq. (32) and Eq. (29), but include the correlated contributions in Eq. (31), Eq. (34) and Eq. (35), i.e., we allow incoherent, optically dark excitonic occupations $N_{1s,\mathbf{Q}}$ to be induced via Coulomb-enhanced Pauli-blocking. Here, the exciton dynamics for the GaAs QW matches remarkably well the electron dynamics in the SBE in Fig. 7(a) with minor deviations at small pulse areas below $1\pi$. In a MoSe2 monolayer, Fig. 7(i), the Rabi flopping dynamics is very similar to the SBE-case in Fig. 7(f) until the pulse maximum is reached, from which on deviations occur, as the coherent excitation of incoherent and optically dark excitonic occupations is less effective compared to the GaAs QW.
@@
In Fig. 7(e,j), we consider all contributions, i.e., we turn on again Eq. (32) as well as Eq. (29) and include Eq. (33), Eq. (43) and Eq. (44). Here, the Rabi-flopping dynamics for a GaAs QW are slightly suppressed compared to Fig. 7(d) due to the excitation-induced dephasing, which occurs via the two-exciton continuum $B_{+,\zeta}$, driven nonlinearly via $P_{1s}P_{1s}$, cf. Eq. (29), as well as via the exciton-two-exciton continuum $Z_{+,\zeta}$ driven by $P_{1s}N_{1s,\mathbf{Q}}$, cf. Eq. (33). For the MoSe2 monolayer, almost all Rabi flops are suppressed, while only a few in the very beginning around 2-2\,ps remain. In both cases, the GaAs QW and the MoSe2 ML, the inclusion of exciton-exciton interaction enhances the overall optically injected exciton density with the most pronounced contribution in the MoSe2 ML at the largest pulse areas. This amplification is caused by the red-shift contribution causing a more effective optical excitation of excitonic transitions, which, in turn, as the pulse area increases, leads to an enhanced coherent excitation of excitonic occupations in fourth order via $\hbar\Omega|P_{1s}|^{2}P_{1s}$, cf. Eq. (34).
@@
All in all, our excitonic theory reproduces the key physics encoded in the semiconductor Bloch equations, which have been proven to reliably model Rabi flops in ultrafast experiments in the moderate Coulomb-interaction regime (GaAs QW) Schülzgen et al. (1999); Binder et al. (1999); Cundiff et al. (1994), while showing major deviations from the SBEs due to exciton-exciton interaction in the strong Coulomb-interaction regime (MoSe2 ML). Here, in circular excitation, Rabi oscillations in the incoherent exciton density are effectively suppressed, while a few Rabi flops in the coherent exciton density can be eventually observed, if the linewidth does not exceed $2\hbar\gamma_{\text{nrad}}=2\,$meV.
Coherent vs. incoherent exciton density:
Figure 8: Dynamics of the coherent exciton density $\frac{1}{\mathcal{A}}|P|^{2}$ (solid lines) and the incoherent exciton density $\frac{1}{\mathcal{A}}\sum_{\mathbf{Q}}N_{\mathbf{Q}}$ (dotted lines) for the full simulations of a GaAs quantum well and a h-BN-encapsulated MoSe2 monolayer from Fig. 7(e,j).
@@
To examine the different dynamical behavior of the GaAs QW and the MoSe2 ML in more detail, we depict the individual coherent ($|P_{1s}|^{2}$, solid lines) and incoherent ($N_{1s,\mathbf{Q}}$, dotted lines) contributions in Fig. 8. It becomes clear, that the Rabi flopping in a GaAs QW occurs in the incoherent part of the total excitonic density $\frac{1}{\mathcal{A}}\sum_{\mathbf{Q}}N_{1s,\mathbf{Q}}$, while in a MoSe2 monolayer, it occurs in the coherent part of the total exciton density $\frac{1}{\mathcal{A}}|P_{1s}|^{2}$. This is a consequence of the weaker Coulomb interaction in the GaAs QW, which does not significantly suppress fermionic Pauli-blocking effects in Eq. (31) (last line) leading to an efficient formation of incoherent, optically dark excitonic occupations $N_{1s,\mathbf{Q}}$ in fourth order, cf. Eq. (34) and Eq. (35). In a MoSe2 ML, due to the stronger Coulomb interaction, fermionic Pauli-blocking effects are reduced, while, at the same time, the Coulomb-induced field renormalization, i.e., the excitonic Rabi energy, is enhanced. This leads to a regime, where Rabi oscillations preferably occur in the coherent exciton density and don’t translate to the incoherent, optically dark exciton density.
Momentum distribution of optically dark excitons:
@@
In Fig. 9, we depict the momentum distribution of the coherently excited excitonic occupations $N_{1s,\mathbf{Q}}$ for a GaAs QW and a MoSe2 ML. Similar to the fully bosonic theory in Ref. Yang et al. (2004), our theory predicts a coherent excitation of excitonic occupations at center-of-mass momenta well outside the light cone (grey dotted vertical line). At increasing pulse areas, the distribution becomes slightly broader. While the distribution in a GaAs QW is well localized in momentum space with an energy width of approximately 45 K, the distribution in a MoSe2 ML is broad with an energy width of approximately 1500 K (note the different COM-momentum scale). This difference is a direct consequence of the different Coulomb interaction, cf. also the corresponding excitonic wave function $\varphi_{\mathbf{q}}$ (blue solid line): A stronger Coulomb interaction enhances the individual electron and hole collisions due to the fermionic substructure in the exciton gas, leading to broader distributed coherently excited excitonic occupations.
Circular vs. linear excitation:

### v2

@@ IV Excitonic Rabi Oscillations
in units of one over area. Within an expansion in electron-hole operator pairs [58, 53], cf. Eq. (7), the total electron density $N$ can also be expressed as:
@@
To study the exciton-density dynamics in Eq. (47), consider two cases, a MoSe2 ML (strong Coulomb coupling) and a GaAs QW (less strong Coulomb coupling). Both systems are resonantly excited at $\hbar\omega_{\text{P}}=E_{1s}$ using a circularly polarized 2-ps pulse. We solve the excitonic equations of motion for the excitonic transitions $P_{1s}$ and excitonic occupations $N_{1s,\mathbf{Q}}$. The equations of motion for the two-exciton continua $B_{+,\zeta,1s}$ and exciton-two-exciton continua $Z_{+,\zeta,1s,\mathbf{Q}}$ for $\xi=\xi^{\prime}=K,\uparrow$ (MoSe2 ML) or $\xi=\xi^{\prime}=\text{heavy-hole~exciton}$ (GaAs QW) are eliminated via a Markov approximation, i.e., we assume an instantaneous build-up of exciton-exciton interaction and neglect memory effects [5], which is reasonable given the comparably long pulse duration. We take only $\mu=1s$ excitonic transitions and occupations into account and neglect any intraexcitonic coherences $N_{\mu,\nu\neq\mu,\mathbf{Q}}$, but include $1s$, $2s$ and $3s$ excitonic states in the calculation of the symmetric two-exciton continua $B_{+,\zeta}$ and symmetric exciton-two-exciton continua $Z_{+,\zeta}$. Due to the circular excitation, antisymmetric “-”-configurations, which host a bound biexciton/exciton-biexciton, are not induced.
For comparison with the excitonic theory, we also solve the well-established semiconductor Bloch equations in Hartree-Fock limit [80] to obtain the density of uncorrelated electrons in Eq. (46):
@@
All material parameters are listed in Tab. 2. To ensure comparable conditions, we assume equal nonradiative dephasings $\gamma_{\text{nrad}}$ for the MoSe2 ML and GaAs QW and neglect any contributions due to reradiation [67]. In a MoSe2 ML, a total (half) linewidth of 1 meV is not far from the radiatively-limited case, which can be achieved by suppressing inhomogeneous broadening via h-BN encapsulation [113, 24, 1], while in a GaAs QW, a total (half) linewidth of 1 meV is much larger than the radiative linewidth of approximately tens of µeV [103] and should be regarded as an effective linewidth for a fair comparison between the different materials. Incorporating inhomogeneous broadening, as typically causing the linewidth in GaAs QWs at temperatures around 4 K [123], would require an ensemble average.
@@
Effective well width $d_{\text{eff}}$ $1.3d$ [81]
@@
Static dielectric constant of GaAs 12.46 (5 K) [89]
@@
h-BN-encapsulated MoSe2 ML
Monolayer width $d$ 0.6527 nm [73]
@@
Static dielectric constant of bulk MoSe2 $\sqrt{8.2\cdot 17.7}$ [78]
@@
Plasmon energy of MoSe2 $\hbar\omega_{\text{pl}}$ 22 eV [71]
@@
CMR [3])
Static dielectric constant of bulk h-BN 4.8 [77]
Interlayer gap $h$ 0.3 nm [40, 37]
@@
GaAs QW MoSe2 ML
@@
Figure 7: Dynamics of the total electron density NN from Eq. (46) as solution of the semiconductor Bloch equations (SBE) (a,f) compared to the total exciton density NN from Eq. (47) as a solution of the excitonic Bloch equations in the coherent limit (COH) (b,g), with additional exciton-exciton scattering contributions (COH+\,+\,$X$) (c,h) and with additional correlated contributions (COH+\,+\,CORR) (d,i) and with all contributions (COH+\,+\,CORR+\,+\,$X$) (e,j) at various pulse areas $\Theta$, cf. Eq. (27), for a 8-nm GaAs QW (left column) and for a MoSe2 ML (right column). The grey-shaded area depicts the normalized pulse intensity.
@@
In the following, we dissect the corresponding interaction contributions to the exciton dynamics and compare the emerging dynamics of a MoSe2 ML and a GaAs QW for circularly-polarized excitation (GaAs QW vs. MoSe2 ML). Then, we identify the role of the coherent and (optically dark) incoherent exciton density (Coherent vs. incoherent exciton density) and examine the momentum distribution of optically dark excitons (Momentum distribution of optically dark excitons). At last, we compare circular and linear excitation in a MoSe2 ML (Circular vs. linear excitation).
@@
GaAs QW vs. MoSe2 ML: In Fig. 7, we depict the dynamics of the electron density, cf. Eq. (46), by solving the semiconductor Bloch equations (SBEs) in Eq. (48) and the dynamics of the total excitonic density, cf. Eq. (47), as solutions of the full excitonic Bloch equations in Eqs. (22)–(45) within the limits discussed above at various pulse areas, defined in Eq. (27). The left column depicts simulations for a GaAs QW and the right column depicts the simulations for a h-BN-encapsulated MoSe2 ML.
@@
In Fig. 7(a,f), which shows the results from solving the SBEs, Eq. (48), Rabi flops are observed, whose period increases with increasing pulse area. In Fig. 7(a) (GaAs QW), up to four Rabi flops can be observed at a pulse area of $2\pi$, while Fig. 7(f) (MoSe2 ML) shows a multitude of Rabi flops. As Rabi oscillations are enhanced due to the occurrence of Coulomb-induced internal fields, cf. last term in the first line in Eq. (48), and the Coulomb interaction in a MoSe2 ML is much stronger compared to a GaAs QW due to the reduced screening, more Rabi flops occur in a MoSe2 ML.
@@
In Fig. 7(b,g), we depict the coherent (COH) limit of the excitonic Bloch equations, cf. Eq. (22) and Eq. (31), which corresponds to the excitonic description in Ref. [68], which does not include nonradiative dephasing. Here, the exciton density decays within 3 ps, as no long-lived incoherent occupations $N_{1s,\mathbf{Q}}$ are induced. In the GaAs QW, Fig. 7(b), one Rabi flop occurs at the largest pulse area, as the nonradiative dephasing $\hbar\gamma_{\text{nrad}}$ suppresses them. In a MoSe2 ML, Fig. 7(g), Rabi flops occur in a similar number compared to the SBE-case in Fig. 7(f), since the stronger Coulomb-induced field renormalization counterbalances the nonradiative dephasing on short timescales.
@@
In Fig. 7(c,h), we depict the coherent limit with exciton-exciton interaction corresponding to two-exciton continua (COH+\,+\,$X$), cf. Eq. (22), Eq. (31), Eq. (33) and Eq. (30), which corresponds to the model developed in Refs. [59, 60, 57]. In this limit, the two-exciton-induced dephasing, which corresponds to the coupling of excitonic transitions to two-exciton continua (excitation-induced dephasing) totally suppresses the (coherent) Rabi oscillations in the GaAs QW and greatly suppresses them in a MoSe2 ML, but not completely, as two Rabi flops of small amplitude still remain. The overall induced coherent exciton density is larger compared to Fig. 7(b,g), as the exciton-exciton interaction reduces the overall excitation-dependent blue shift due to bandgap reduction and phase-space filling in Eq. (31) (first line) enabling a more efficient optical excitation of electron-hole pairs. This is especially relevant in a MoSe2 ML, as the enhanced Coulomb interaction causes a stronger exciton-exciton interaction and hence a stronger excitation-induced dephasing and -attenuation of the blue shift.
@@
In Fig. 7(d,i), we turn off the exciton-exciton (XX) interaction in Eq. (33) and Eq. (30), but include the correlated contributions in Eq. (32), Eq. (35) and Eq. (36), i.e., we allow incoherent, optically dark excitonic occupations $N_{1s,\mathbf{Q}}$ to be induced via Coulomb-enhanced Pauli-blocking. Here, the exciton dynamics for the GaAs QW matches remarkably well the electron dynamics in the SBE in Fig. 7(a) with minor deviations at small pulse areas below $1\pi$. In a MoSe2 ML, Fig. 7(i), the Rabi flopping dynamics is very similar to the SBE-case in Fig. 7(f) until the pulse maximum is reached, from which on deviations occur, as the coherent excitation of incoherent and optically dark excitonic occupations is less effective compared to the GaAs QW.
@@
In Fig. 7(e,j), we consider all contributions, i.e., we turn on again Eq. (33) as well as Eq. (30) and include Eq. (34), Eq. (44) and Eq. (45). Here, the Rabi-flopping dynamics for a GaAs QW are slightly suppressed compared to Fig. 7(d) due to the excitation-induced dephasing, which occurs via the two-exciton continuum $B_{+,\zeta}$, driven nonlinearly via $P_{1s}P_{1s}$, cf. Eq. (30), as well as via the exciton-two-exciton continuum $Z_{+,\zeta}$ driven by $P_{1s}N_{1s,\mathbf{Q}}$, cf. Eq. (34). For the MoSe2 ML, almost all Rabi flops are suppressed, while only a few in the very beginning around 2-2\,ps remain. In both cases, the GaAs QW and the MoSe2 ML, the inclusion of exciton-exciton interaction enhances the overall optically injected exciton density with the most pronounced contribution in the MoSe2 ML at the largest pulse areas. This amplification is caused by the red-shift contribution causing a more effective optical excitation of excitonic transitions, which, in turn, as the pulse area increases, leads to an enhanced coherent excitation of excitonic occupations in fourth order via $\hbar\Omega|P_{1s}|^{2}P_{1s}$, cf. Eq. (35).
We note that, in general, excitonic Rabi oscillations in semiconductors are – in contrast to an undamped ideal two-level system, cf. Fig. 8(a) – never fully modulated, even if we neglect any damping, cf. Fig. 8(b). This behavior already occurs without any Coulomb interaction and is due to the fact, that exciting a semiconductor at the band gap with $\mathbf{k}=\mathbf{0}$ creates a multitude of electron-hole pairs, i.e., an ensemble of two-level systems, which experience a slightly different Rabi energy each with increasing difference at increasing quasi-momentum $\mathbf{k}$. The Rabi-flopping dynamics of the resulting total electron density are then governed by the average over many two-level systems oscillating with slightly different Rabi frequencies causing a superposition resulting in a partial modulation of the oscillations only. In case of nonzero damping, cf. Fig. 8(c,d), Rabi oscillations in a semiconductor without Coulomb interaction are even more damped. In general, the Coulomb interaction not only retrieves the Rabi flops but also enhances their number by a factor of approximately two [107, 19], cf. Fig. 7, due to excitonic effects on Hartree-Fock level. However, a full modulation of Rabi-oscillations due to two-level-like Pauli-blocking – even for strict resonant excitation of the exciton – cannot occur, since exciton-exciton interactions obscure the Pauli-blocking property.
Occupation (1)
Single TLS SBE without Coulomb
Density ($10^{10}\,$cm-2)
Figure 8: Dynamics of the excited occupation of single two-level system as a solution of the single-state limit of the SBE without Coulomb interaction, the optical Bloch equations [2] (a,c) and dynamics of the electron density NN from Eq. (46) as a solution of the SBE without Coulomb interaction (b,d) for a 8-nm GaAs QW for various pulse areas. The grey-shaded area depicts the normalized pulse intensity.
@@
All in all, our excitonic theory reproduces the key physics encoded in the semiconductor Bloch equations, which have been proven to reliably model Rabi flops in ultrafast experiments in the moderate Coulomb-interaction regime (GaAs QW) [107, 19, 29], while showing major deviations from the SBEs due to exciton-exciton interaction in the strong Coulomb-interaction regime (MoSe2 ML). Here, in circular excitation, Rabi oscillations in the incoherent exciton density are effectively suppressed, while a few Rabi flops in the coherent exciton density can be eventually observed, if the linewidth does not exceed $2\hbar\gamma_{\text{nrad}}=2\,$meV.
@@
Coherent vs. incoherent exciton density: To examine the different dynamical behavior of the GaAs QW and the MoSe2 ML in more detail, we depict the individual coherent ($|P_{1s}|^{2}$, solid lines) and incoherent ($N_{1s,\mathbf{Q}}$, dotted lines) contributions in Fig. 9. It becomes clear, that the Rabi flopping in a GaAs QW occurs in the incoherent part of the total excitonic density $\frac{1}{\mathcal{A}}\sum_{\mathbf{Q}}N_{1s,\mathbf{Q}}$, while in a MoSe2 ML, it occurs in the coherent part of the total exciton density $\frac{1}{\mathcal{A}}|P_{1s}|^{2}$. This is a consequence of the weaker Coulomb interaction in the GaAs QW, which does not significantly suppress fermionic Pauli-blocking effects in Eq. (32) (last line) leading to an efficient formation of incoherent, optically dark excitonic occupations $N_{1s,\mathbf{Q}}$ in fourth order, cf. Eq. (35) and Eq. (36). In a MoSe2 ML, due to the stronger Coulomb interaction, fermionic Pauli-blocking effects are reduced, while, at the same time, the Coulomb-induced field renormalization, i.e., the excitonic Rabi energy, is enhanced. This leads to a regime, where Rabi oscillations preferably occur in the coherent exciton density and don’t translate to the incoherent, optically dark exciton density.
Figure 9: Dynamics of the coherent exciton density $\frac{1}{\mathcal{A}}|P|^{2}$ (solid lines) and the incoherent exciton density $\frac{1}{\mathcal{A}}\sum_{\mathbf{Q}}N_{\mathbf{Q}}$ (dotted lines) for the full simulations of a GaAs QW and a h-BN-encapsulated MoSe2 ML from Fig. 7(e,j).
Momentum distribution of optically dark excitons: In Fig. 10, we depict the momentum distribution of the coherently excited excitonic occupations $N_{1s,\mathbf{Q}}$ for a GaAs QW and a MoSe2 ML. Similar to the fully bosonic theory in Ref. [133], our theory predicts a coherent excitation of excitonic occupations at center-of-mass momenta well outside the light cone (grey dotted vertical line). At increasing pulse areas, the distribution becomes slightly broader. While the distribution in a GaAs QW is well localized in momentum space with an energy width of approximately 45 K, the distribution in a MoSe2 ML is broad with an energy width of approximately 1500 K (note the different COM-momentum scale). This difference is a direct consequence of the different Coulomb interaction, cf. also the corresponding excitonic wave function $\varphi_{\mathbf{q}}$ (blue solid line): A stronger Coulomb interaction enhances the individual electron and hole collisions due to the fermionic substructure in the exciton gas, leading to broader distributed coherently excited excitonic occupations.
@@
Circular vs. linear excitation: In Fig. 11, we depict the dynamics of the total exciton density at the $K$ valley from Eq. (47), i.e., $\xi=\xi^{\prime}=K,\uparrow$, for circular (a), cf. also Fig. 7(j), and linear (b) excitation. For the linear-excitation case, we include coherent transitions at the $K$ and $K^{\prime}$ valleys $P{\vphantom{P}}_{1s}^{K}$, $P{\vphantom{P}}_{1s}^{K^{\prime}}$ as well as intravalley incoherent occupations $N\vphantom{N}_{1s,\mathbf{Q}}^{K,K}$, $N\vphantom{N}_{1s,\mathbf{Q}}^{K^{\prime},K^{\prime}}$ and neglect any intervalley occupations $N\vphantom{N}_{1s,\mathbf{Q}}^{K,K^{\prime}}$$N\vphantom{N}_{1s,\mathbf{Q}}^{K^{\prime},K}$. While in circular excitation, some Rabi flops of the coherent exciton density remain, cf. Fig. 9 and the discussion above, in linear excitation, even the Rabi flops in the coherent exciton density are almost totally suppressed. The reason is the larger two-exciton/exciton-two-exciton phase space: In circular excitation, only symmetric (++) intravalley two-exciton continua $B\vphantom{B}_{+,\zeta}^{K,K,K,K}$ and exciton-two-exciton continua $Z\vphantom{Z}_{+,\zeta,\mathbf{Q}}^{K,K,K,K,K,K}$ are induced. In linear excitation, the $K$ valley and the $K^{\prime}$ valley are equally excited, so that additional symmetric (++) and antisymmetric (-) intervalley two-exciton continua $B\vphantom{B}_{\pm,\zeta}^{K,K^{\prime},K^{\prime},K}$$B\vphantom{B}_{\pm,\zeta}^{K^{\prime},K,K,K^{\prime}}$ and exciton-two-exciton continua $Z\vphantom{Z}_{\pm,\zeta,\mathbf{Q}}^{K^{\prime},K^{\prime},K,K,K^{\prime},K^{\prime}}$$Z\vphantom{Z}_{\pm,\zeta,\mathbf{Q}}^{K,K,K^{\prime},K^{\prime},K,K}$ are induced. These yield a dynamic coupling between the excitonic transitions at the $K$ valley $P{\vphantom{P}}_{1s}^{K}$ and at the $K^{\prime}$ valley $P{\vphantom{P}}_{1s}^{K^{\prime}}$, hence further increasing the excitation-induced dephasing. Moreover, the formation of a bound ($b$) biexciton $B\vphantom{B}_{-,\zeta=b}^{K,K^{\prime},K^{\prime},K}$$B\vphantom{B}_{-,\zeta=b}^{K^{\prime},K,K,K^{\prime}}$ and bound exciton-biexciton $Z\vphantom{Z}_{-,\zeta=b,\mathbf{Q}}^{K^{\prime},K^{\prime},K,K,K^{\prime},K^{\prime}}$$Z\vphantom{Z}_{-,\zeta=b,\mathbf{Q}}^{K,K,K^{\prime},K^{\prime},K,K}$ becomes possible, further enhancing the excitation-induced dephasing. Hence, due to the overall larger number of Coulomb-interaction-induced exciton-exciton correlations and the resulting dynamic coupling of the excitonic transitions $P$ between the $K$ and $K^{\prime}$ valley, the excitation-induced dephasing is larger in linear excitation compared to circular excitation causing a suppression of the Rabi oscillations in the coherent exciton density. This theoretical finding is consistent with recent experiments in a MoSe2 ML, where a clear Rabi splitting due to an occupation grating, which is insensible to coherent excitation-induced dephasing, has been measured, but no Rabi oscillations could be observed [103].